Suppose you found $1,000 hidden in your mattress and deposited it in a demand deposit account at your bank. If the reserve requirement was 20 percent, the deposit would directly create in excess reserves and ultimately lead to a total increase in the money supply, if all banks in the system lend out 100 percent of their excess reserves.


A. $800 ; $4,000

B. $800 ; $5,000

C. $1,000 ; $4,000

D. $1,000 ; $5,000
Answer #1

Deposit in demand deposit = $1,000

Reserve Requirement = 20%

Calculate the money multiplier

Money multiplier = 1 ÷ Reserve Requirement

Money multiplier = 1 ÷ 20% = 5

Excess reserve will be = Deposit – Reserve

Excess reserve will be = $1,000 – 20% of $1,000 = $800

Increase in money supply = Excess Reserve * Money Multiplier

Increase in money supply = Excess Reserve * Money Multiplier

Increase in money supply = $800* 5 = $4,000

Answer – A. $800; $4,000